# Modelling Lateral Spread in Wire Flat Rolling

### v1

@@ Abstract
A mathematical model for wire rolling is presented, focusing on predicting the lateral spread. This provides, for the first time, an analytic model of lateral spread without any fitting parameters. The model is derived directly from the governing equations; in order to obtain a simple model, a rigid perfectly plastic material is assumed and simplified with the assumption of small thickness and width compared to the length of the roll gap (effectively a thin-wire large-roller assumption). The mathematical model is compared against experiments performed on stainless steel wire using a $100\,\mathrm{mm}$ diameter roller, and is shown to provide accurate predictions of lateral spread across a wide range of wire diameters ($2.96\,\mathrm{mm}$$7.96\,\mathrm{mm}$) and reduction ratios (20%–60%), all without the need for fitting parameters. The model requires only seconds to compute and serves as a robust tool for validating FE results, guiding process design, and laying the foundation for future improved models.
@@
journal: J. Matter. Process. Technol.
@@ 1 Introduction
Applications including sawblades, springs, piston rings, and transformers depend critically on flattened wires [1]. These wires are usually made by a flat rolling technique where a wire with a circular cross-section is cold rolled between cylindrical rolls — sometimes through several passes — to achieve a particular width and thickness (figure 1).
Figure 1: A diagram of wire flat rolling process; the wire initially has a circular cross-section with diameter $d_{0}$ and is flattened to a barrel shape cross section with the lateral spread $W_{t}$ and the contact width $W_{c}$. This is a modified version of Figure 1 from Carlsson [2].
@@
In cold rolling of a sheet, deformation in the lateral direction often remains within the elastic range due to the product’s geometry. However, this is not the case when an initial width-to-thickness ratio ($W_{0}/2h_{0}$) is less than 6, such as in wide strip or plate, or as low as 1 in round wire [3]. In such cases, the plastic flow in the roll gap at the start of rolling schedule is inherently three-dimensional, complicating the analysis. For round wire, this complexity is further increased by the transformation of a round cross-section into a rectangular shape with bulged edges during the first pass.
@@
Although research exists on the modelling of lateral spread in thick plate rolling [3, 4, 5, 6, 7], wire flat rolling has received less attention. Kazeminezhad and Taheri published a series of papers studying different parameters in wire rolling [8, 9, 10]. Initially, Kazeminezhad and Taheri [8] developed a relationship for the width of the contact area between the rolls and the wire. Their equation was based on experimental and FE observations of the formation of X-shaped shear bands in the wire cross-section during the rolling process [11, 12], along with the assumption that as the height reduction increases, the shear bands rotate while maintaining a constant length. This equation was then applied alongside the slab method [9], leading to a pressure hill distribution. The study employs the Tresca yield criterion, assuming the longitudinal stress $\sigma_{xx}$ to be the minimum stress, the vertical stress $\sigma_{yy}$ to be the maximum stress, and the compressive stress in the lateral direction, $\sigma_{zz}$, to lie between them. Although $\sigma_{zz}$ is implicitly included in the yield function, and therefore in calculating the roll pressure, since the value of $\sigma_{zz}$ is unknown, it must be approximated to find the lateral spread. Kazeminezhad and Taheri [8] initially assumed $\sigma_{zz}$ to take a value between plane strain and plane stress, written as $m\sigma_{yy}$, where $m$ is a fitting parameter taking values between 0 and 0.5. Using flow rule equations and setting $\sigma_{xx}$ to be zero at the end of the roll gap, they then wrote
$\frac{d\varepsilon_{zz}}{-d\varepsilon_{yy}}=\frac{\ln\frac{W_{t}}{d_{0}}}{\ln\frac{d_{0}}{2h_{1}}}=\frac{1-2m}{2-m},$ (1)
@@
where $2h_{1}$ is the final height, $d_{0}$ is the diameter of the wire, and $W_{t}$ is the lateral spread of the wire. The first use of this formulation has been attributed to Hill [13] for plate rolling and has also been derived by several other authors [e.g. 3, 14], who experimentally observed a linear relationship between $\ln(W_{t}/W_{0})$ and $\ln(h_{0}/h_{1})$. The factor $m$ and another factor of linearity must then be found empirically from experiments with a proviso that each resulting equation only gives reasonable lateral spread prediction within the ranges of conditions for which they were empirically determined [3]. For example, for low and high carbon steel Kazeminezhad and Taheri [8] wrote equation (1) as
@@
Among a wide range of empirical formulations suggested for predicting lateral spread [3], the one proposed by Kobayashi [15] is often referred to as being reliable for round wire[10, 1, 16]. Vallellano et al. [16] and Utsunomiya et al. [1] attribute the following equation to Kobayashi [15], and we will refer to it here by this name:
@@
Similar to other metal-forming processes, finite element (FE) analysis has provided valuable insights into wire rolling. Vallellano et al. [16] performed a 3D numerical analysis using the Abaqus finite element software to study contact stress distributions, residual stresses, and lateral spread. While 3D FE analysis offers highly detailed information, it is computationally expensive and requires validation against a reliable reference. Whether the objective is process design or FE validation, a review of the literature further reveals a notable gap in the availability of models with well-defined, traceable assumptions for the rolling of round wire.
Here, a mathematical model is developed with the assumption of a small friction coefficient and large length-to-thickness ratio, consistent with the slab method of Kazeminezhad and Taheri [9]. The further assumption of comparably small width is made, which simplifies the problem to plane-stress, allowing for finding the lateral spread without the need for fitting parameters or solving a problem in a complicated 3D stress state. Even so, the derivation of equations in all directions and the use of asymptotic analysis provide a foundation for further improvements to the model in the future, unlike the slab method, which remains inherently limited.
The mathematical model is explained in Section 2. The non-dimensionalised governing equations in all directions are derived in Section 2.1 for a rigid perfectly plastic material, exploiting small parameters of the ratio of half-thickness and width to length of roll gap and the friction coefficient. These assumptions reduce the general deformation state to plane-stress deformation at the leading order, which is solved in Section 2.2. Section 2.3 summarises the solutions and outlines the computational methods. The experimental procedure for obtaining data is detailed in Section 3, which is used to validate lateral spread predictions in Section 4.1, alongside FE data from Vallellano et al. [16]. Similarly, the roll pressure predictions are validated against FE data from Vallellano et al. [16] and presented in Section 4.2. Finally, a discussion on potential directions for future research is provided in Section 5.
@@ 2 Mathematical model
The schematic diagram of the model is shown in figure 2. In the first pass of rolling, the round wire undergoes deformation, resulting in a barrel-like cross-section. With the goal of the simplest possible model a set of assumptions is made to evade the complexity arising from round cross section at the entrance and barrel-like cross-section afterwards. Firstly, instead of accounting for the circular cross-section at the entrance, it is assumed that at some point after entry, the initially circular cross-section has transformed into a rectangular shape while preserving its area. This stage is chosen as the beginning of the roll gap in the model and is marked A, in figure 2.
@@
Therefore, we may write
$4\hat{h}_{\text{A}}^{2}$ $=\frac{\pi}{4}{\hat{d}_{0}}^{2}$$\Rightarrow$ $2\hat{h}_{\text{A}}$ $=\frac{\sqrt{\pi}}{2}\hat{d}_{0},$ (4)
where $\hat{h}_{\text{A}}$ is the half thickness of the wire corresponding to location A. As the quantities will later be non-dimensionalised, and to clearly distinguish between dimensional and non-dimensional parameters, we adopt the convention that, for the remainder of the paper, variables with hats denote dimensional quantities, while unhatted variables represent their dimensionless counterparts. Secondly, as the rolling progresses, a square cross-section is assumed to progressively flatten into a rectangular shape, the area of which is assumed to be the same as the area of the real cross-section with bulged edges (see figure 2). Thirdly, the bulged edges are approximated as a half-circle, implying that the bulged radius is taken as half the thickness. This allows the model with a rectangular cross-section to be correlated to the real process. As will be shown later, this set of assumptions affects the pressure prediction but not the lateral spread prediction.
In the following, the governing equations are written for rigid perfectly plastic material rolled with rigid rolls while Coulomb friction is imposed between the rolls and material with a constant friction coefficient.
2.1 Scaling and non-dimensionalising the equations
The exploitation of small dimensionless parameters is an important part of the analysis here. We therefore rescale the variables in terms of relevant dimensional scales. The horizontal distance is scaled with the characteristic roll-gap length, $\hat{\ell}$, such that $\hat{x}=\hat{\ell}x$, and the vertical distance with the initial sheet half-thickness, $\hat{h}_{\text{A}}$, so that $\hat{y}=\hat{h}_{\mathrm{A}}y$. For circular rolls of radius $\hat{R}$, the roll-gap length $\hat{\ell}$ is equal to
^$\hat{\ell}=\sqrt{2\hat{R}(\hat{h}_{\text{A}}-\hat{h}_{1})-(\hat{h}_{\text{A}}-\hat{h}_{1})^{2}},$ (5)
@@ 2.1 Scaling and non-dimensionalising the equations
The velocity $(\hat{u},\hat{v},\hat{w})$, is scaled with the wire velocity at point A and to balance incompressibility as $(\hat{R}\hat{U}_{\text{A}},\delta\hat{R}\hat{U}_{\text{A}},\delta\hat{R}\hat{U}_{\text{A}})$. It should be noted that the sheet velocity at any point, including $U_{\text{A}}$, is initially unknown and must be determined based on the roll velocities after finding the location of the neutral point.
@@
Hydrostatic pressure, $\hat{p}$, and the normal Cauchy stresses, $\hat{\sigma}_{xx}$ and $\hat{\sigma}_{yy}$, and their deviatoric components, $\hat{s}_{xx}$ and $\hat{s}_{yy}$, are all scaled with shear yield stress, $\hat{\kappa}$. The assumption of a small friction coefficient is encoded by setting $\beta=\delta\mu$, where $\beta$ is a quantity of approximate unit magnitude. Similarly, the shear stress, $\hat{\sigma}_{xy}$ is scaled with $\delta\hat{\kappa}$ resulting from small friction. To determine the order of magnitude of stress components in the $z$-direction the boundary condition is analysed. The lateral edges of the wire must be tension-free so that the material can move in this direction ($z$-direction). This condition may be written as $\boldsymbol{\sigma\cdot s}=0$, where $\boldsymbol{s}$ is the unit vector normal to the wire lateral edges. For a rectangular cross section $\boldsymbol{s}=(-\delta\mathrm{d}b/\mathrm{d}x,0,1)$. Therefore, at $z=b(x)$, we get
@@
$\delta\frac{\mathrm{d}b}{\mathrm{d}x}\sigma_{xx}$ $=\hat{\sigma}_{xz}/\hat{\kappa},$ (6a)
$\delta^{2}\frac{\mathrm{d}b}{\mathrm{d}x}\sigma_{xy}$ $=\hat{\sigma}_{yz}/\hat{\kappa},$ (6b)
$\delta\frac{\mathrm{d}b}{\mathrm{d}x}\hat{\sigma}_{xz}/\hat{\kappa}$ $=\hat{\sigma}_{zz}/\hat{\kappa},$ (6c)
where $\mathrm{d}b/\mathrm{d}x$, $\sigma_{xx}$, and $\sigma_{xy}$, whose orders of magnitude are preserved, are written in dimensionless form, and $\hat{\sigma}_{zz}$, $\hat{\sigma}_{xz}$, and $\hat{\sigma}_{yz}$, whose orders of magnitude are unknown, are written in dimensional form. The set of equations 6 implies that $\hat{\sigma}_{xz}$ is $O(\delta)$ and the normal stress in the lateral direction, $\hat{\sigma}_{zz}$ and the shear stress component $\hat{\sigma}_{yz}$ are $O(\delta^{2})$. Although this holds at the lateral edges, we extend the assumption to apply these orders of magnitude throughout the modelling region. Consequently, $\hat{\sigma}_{zz}$ and $\hat{\sigma}_{yz}$ are scaled with $\delta^{2}\hat{\kappa}$, while $\hat{\sigma}_{xz}$ is scaled with $\delta\hat{\kappa}$. As will be shown later, this approach leads to the simplest yet consistent solution, where the horizontal velocity and normal stresses depend only on $x$. To complete the non-dimensionalisation, the plastic parameter, $\hat{\lambda}$, is scaled with $\hat{U}_{A}/\hat{\kappa}\hat{\ell}$.
Under the scaling introduced, the dimensionless form of force balance in each direction is
@@
The dimensionless form of the incompressibility and flow rule relations is
$\frac{\partial{u}}{\partial{x}}+\frac{\partial{w}}{\partial{z}}+\frac{\partial{v}}{\partial{y}}=0,$ (8a)
@@
$\frac{\partial{u}}{\partial{y}}+\delta^{2}\frac{\partial{v}}{\partial{x}}=\delta^{2}\frac{1}{2}\lambda\sigma_{xy},$ (8e)
$\frac{\partial{u}}{\partial{z}}+\delta^{2}\frac{\partial{w}}{\partial{x}}=\delta^{2}\frac{1}{2}\lambda\sigma_{xz},$ (8f)
$\frac{\partial{v}}{\partial{z}}+\frac{\partial{w}}{\partial{y}}=\delta^{2}\frac{1}{2}\lambda\sigma_{yz}.$ (8g)

### v2

@@ Abstract
A mathematical model for wire rolling is developed, focusing on predicting the lateral spread. This provides, for the first time, an analytic model of lateral spread without any fitting parameters. The model is derived directly from the governing equations, assuming a rigid, perfectly plastic material and exploiting the thinness of the wire (in thickness and width) relative to the roller size. Results are compared against experiments performed on stainless steel wire using $100\,\mathrm{mm}$ diameter rolls, demonstrating accurate predictions of lateral spread across a wide range of wire diameters ($2.96\,\mathrm{mm}$$7.96\,\mathrm{mm}$) and reduction ratios (20%–60%), all without the need for fitting parameters. Since the model requires only seconds to compute, the model’s valid range is explored for varying roll diameter, wire diameter, and reduction ratio, and their effects on the resulting lateral spread characterized. The model can serve as a robust tool for validating FE results, guiding process design, and laying the foundation for future improved models. Matlab code to evaluate the model is provided in the supplementary material.
@@
journal: Eur. J. Mech. A/Solids
@@ 1 Introduction
Applications including sawblades, springs, piston rings, and transformers depend critically on flattened wires (utsunomiya2001three). These wires are usually made by a flat rolling technique, where a wire with a circular cross-section is cold rolled between cylindrical rolls — sometimes through several passes — to achieve a particular width and thickness (as illustrated in figure 1).
Figure 1: A diagram of wire flat rolling process; the wire initially has a circular cross-section with diameter $d_{0}$ and is flattened to a barrel shape cross section with the lateral spread $W_{t}$ and the contact width $W_{c}$. Adapted from Figure 1 of carlsson1998contact.
@@
In cold rolling of a sheet, deformation in the lateral direction often remains within the elastic range due to the product’s geometry. However, this is not the case when an initial width-to-thickness ratio is less than 6, such as in wide strip or plate, or as low as 1 in round wire (chitkara1966some). In such cases, the plastic flow in the roll gap at the start of rolling schedule is inherently three-dimensional, complicating the analysis. For round wire, this complexity is further increased by the transformation of a round cross-section into a rectangular shape with bulged edges during the first pass.
@@
Although research exists on the modelling of lateral spread in thick plate rolling (e.g. chitkara1966some; lahoti1974hill; oh1975approximate; Nikhaily; Kennedy), wire flat rolling has received less attention. kazeminezhad2008error (kazeminezhad2005width; kazeminezhad2006pressure; kazeminezhad2008error) published a series of papers studying different parameters in wire rolling. kazeminezhad2005width developed a relationship for the width of the contact area between the rolls and the wire. Their equation was based on experimental and FE observations of the formation of X-shaped shear bands in the wire cross-section during the rolling process (Semiatin1984FormabilityAW; pesin2002mathematical), along with the assumption that as the height reduction increases, the shear bands rotate while maintaining a constant length. kazeminezhad2006pressure then applied this equation alongside the slab method, leading to a pressure hill distribution. The study employs the Tresca yield criterion, assuming the longitudinal stress $\sigma_{xx}$ to be the minimum stress, the vertical stress $\sigma_{yy}$ to be the maximum stress, and the compressive stress in the lateral direction, $\sigma_{zz}$, to lie between them. Although $\sigma_{zz}$ is implicitly included in the yield function, and therefore in calculating the roll pressure, the value of $\sigma_{zz}$ is unknown and must be approximated to find the lateral spread. kazeminezhad2005width therefore assumed that $\sigma_{zz}$ takes a value between plane strain and plane stress, written as $\alpha\sigma_{yy}$, where $\alpha$ is a fitting parameter taking values between 0 and 0.5. Using flow rule equations and setting $\sigma_{xx}$ to be zero at the end of the roll gap, they then wrote
$\frac{d\varepsilon_{zz}}{-d\varepsilon_{yy}}=\frac{\ln\frac{W_{t}}{d_{0}}}{\ln\frac{d_{0}}{2h_{1}}}=\frac{1-2\alpha}{2-\alpha},$ (1)
@@
where $W_{t}$ is the lateral spread, $d_{0}$ is the initial diameter, and $2h_{1}$ is the final height of the wire. The first use of this formulation has been attributed to hill1955 for plate rolling and has also been derived by several other authors (e.g. chitkara1966some; applied1961formula), who experimentally observed a linear relationship between $\ln(W_{t}/W_{0})$ and $\ln(h_{0}/h_{1})$, where $W_{0}$ and $h_{0}$ are initial half width and half height of the slab, respectively. The factor $\alpha$ and another factor of linearity must then be found empirically from experiments, with a proviso that each resulting equation only gives reasonable lateral spread prediction within the ranges of conditions for which they were empirically determined (chitkara1966some). For example, for low and high carbon steel kazeminezhad2005width wrote equation (1) as
@@
Among a wide range of empirical formulations suggested for predicting lateral spread (chitkara1966some), the one proposed by kobayashi1978influence is often referred to as being reliable for round wire (utsunomiya2001three; kazeminezhad2008error; vallellano2008analysis). utsunomiya2001three and vallellano2008analysis attribute the following equation to kobayashi1978influence, and we will refer to it here by this name:
@@
Finite element (FE) analysis has provided valuable insights into wire rolling, similar to other metal-forming processes. vallellano2008analysis performed a 3D numerical analysis using the Abaqus finite element software to study contact stress distributions, residual stresses, and lateral spread. While 3D FE analysis offers highly detailed information, it is computationally expensive, especially when it is used in design iterations where several simulations are required. Although some studies have investigated FE simulations of wire rolling (carlsson1998contact; vallellano2008analysis; hwang2021influence), computational cost data are generally not reported. However, to give an indication, a 2D FE simulation of half-of-the-thickness sheet rolling required 15.9 CPU hours for a reliable case (30 elements through the half-thickness) and 0.25 CPU hours for a coarse case (5 elements) (flanagan2025); since 3D wire-rolling simulations are considerably more complex, they are expected to be significantly more expensive yet. In addition to high computational cost, FE results require validation against a reliable reference. Whether the objective is process design or FE validation, a review of the literature further reveals a notable gap in the availability of models with well-defined, traceable assumptions for the rolling of round wire.
Using similar techniques to the asymptotic mathematical modelling of sheet rolling (erfanian2025), here an asymptotic mathematical model is developed for the regime of a small friction coefficient and large length-to-thickness ratio, consistent with the slab method of kazeminezhad2006pressure. The further assumption that the width is comparably smaller than the length of wire in the roll gap simplifies the problem to plane-stress, where the stress in $z$ direction is negligible. This allows for finding the lateral spread without the need for fitting parameters or solving a problem in a complicated 3D stress state. Even so, the derivation of equations in all directions and the use of asymptotic analysis provide a foundation for further improvements to the model in the future, unlike the slab method, which remains inherently limited.
An outline of the paper is as follows. The mathematical model is explained in Section 2, beginning with section 2.1 introducing the simplifying assumptions. The scaling parameters and resulting non-dimensionalised governing equations are given in Section 2.2 for a rigid perfectly plastic material, and their solution is derived in Section 2.3. Section 2.4 summarises the solutions and outlines the computational methods used to evaluate them. The experimental procedure for obtaining data is detailed in Section 3, which is used to validate the lateral spread predictions of the model in Section 4.1, alongside FE data from vallellano2008analysis. The Longitudinal spread is compared against experimental data in Section 4.2 and the roll pressure predictions are validated against FE data from vallellano2008analysis and presented in Section 4.3. Having validated the model, a parametric study is presented in Section 5, exploring both the range of validity of the model in Section 5.1 and the influence of parameters including friction, reduction, and roll diameter on the lateral spread in Section 5.2. Finally, a discussion and potential directions for future research are provided in Section 6.
@@ 2 Mathematical model
2.1 Assumptions
We first make two assumptions to help in modelling practical wire rolling: section 2.1.1 concerns the initial transformation of a cylindrical wire in the first roll-stand, and section 2.1.2 concerns the bulged edges of the subsequently rolled wire. These assumptions allow for a simpler model based on rectangular cross-sections to be developed in section 2.2.
2.1.1 Circular to rectangular deformation
A schematic diagram of the model is shown in figure 2.
@@ 2.1.1 Circular to rectangular deformation
In the first pass of rolling, the round wire undergoes deformation. Here, we assume that at some point after entry, the initially circular cross-section has transformed into a rectangular shape without changing the cross-sectional area. This stage is chosen as the beginning of the roll gap in the asymptotic model, and is marked A in figure 2. Therefore, instead of solving the equation from the actual roll-gap entrance, where the cross-section is circular, the information is translated to point A in the roll gap, and the equations are solved from this point onwards. The intuition behind this assumption is that we hypothesise it is much easier for a circular cross-section wire at the beginning of the process — where only a small region is in contact with the rolls — to deform into a rectangular shape with a flat top than to elongate. It should also be noted that the assumption of an unchanged cross-sectional area between the entrance and point A does not imply that the wire length remains constant, as no assumption about the cross-sectional area is made beyond point A. All of these assumptions, although physically justifiable, remain approximations that are validated in section 4 through comparison with experimental results. The location A is determined by equating the cross-sectional area at this location to that of the initial circular wire. Thus, if $\hat{h}_{\!\>\!\mathrm{A}}$ and $\hat{b}_{\!\>\!\mathrm{A}}$ are the half-thickness and half-width of the wire at location A, and we assume that the initial rectangular cross-section has aspect ratio $\hat{b}_{\!\>\!\mathrm{A}}/\hat{h}_{\!\>\!\mathrm{A}}=a$, we may write
$4\hat{h}_{\!\>\!\mathrm{A}}\hat{b}_{\!\>\!\mathrm{A}}$ $=4a\hat{h}_{\!\>\!\mathrm{A}}^{2}=\frac{\pi}{4}{\hat{d}_{0}}^{2}$ $\Rightarrow$ $\hat{h}_{\!\>\!\mathrm{A}}$ $=\sqrt{\frac{\pi}{16a}}\hat{d}_{0}=\frac{\hat{b}_{\!\>\!\mathrm{A}}}{a}.$ (4)
In order to preserve the symmery of the initially cylindrical wire, we will assume that $a=1$ throughout this work unless explicitly stated; this assumption is further justified in A by comparison with experimental results. As the quantities will later be non-dimensionalised, and to clearly distinguish between dimensional and non-dimensional parameters, we adopt the convention that, for the remainder of the paper, variables with hats denote dimensional quantities, while unhatted variables represent their dimensionless counterparts.
Note that this circular-to-square transition is only relevant to the first rolling pass, and for subsequent passes point A is at the roll-gap entrance, $\hat{h}_{\!\>\!\mathrm{A}}$ and $\hat{b}_{\!\>\!\mathrm{A}}$ are the half-thickness and half-width of the wire at the entrance, and equation (4) is not needed.
2.1.2 Bulged edges of rectangular wire
In practice, the wire never has a perfectly rectangular cross-section but has bulged edges (see figure 2). A second assumption is therefore that, as the rolling progresses, a square cross-section is progressively flattened into a rectangular shape, and the bulged edges are approximated as half-circles whilst maintaining the same area as the perfect rectangle (also shown in figure 2). Consequently,
$2\hat{W}_{c}\hat{h}+\pi\hat{h}^{2}$ $=4\hat{b}\hat{h} \displaystyle\Rightarrow$ $\hat{W}_{c}$ $=2\hat{b}-\pi\hat{h}/2,$ (5)
where $\hat{W}_{c}$ is the width in contact with the roll, and the total lateral spread $\hat{W}_{t}$ is given by
$\hat{W}_{t}=\hat{W}_{c}+2\hat{h}$. (6)
This allows a model of the deformation of a rectangular cross-section to be correlated with the real process with a bulged cross-section. As will be shown later, this set of assumptions allows for the simplest possible model, while having minimal effect on lateral spread prediction.
In the following, we rigorously derive a leading-order asymptotic model for the rolling of a rectangular wire. The ad hoc model previously introduced for the initial transformation of a circular wire into a square cross-section enables this asymptotic formulation to be applied to circular wires as well, which will be validated a posteriori through comparison with FE results. The governing equations are written for rigid perfectly plastic material rolled with rigid rolls while Coulomb friction is imposed between the rolls and material with a constant friction coefficient. The equations are formulated for a regime characterised by a small friction coefficient between the roll and the wire, as well as a wire width and thickness much smaller than its length within the roll gap. The latter condition effectively corresponds to a small wire diameter compared to the roll diameter.
2.2 Scalings and the non-dimensionalised governing equations
Asymptotic (or perturbation) methods (hinch) are widely used in applied mechanics, exploiting the smallness of an inherent parameter to obtain approximate analytical/semi-analytical solutions with controllable errors. The first step in applying this technique is to rescale the governing equations and parameters based on the inherent scales of the problem. This process removes physical units and expresses all quantities in a comparable dimensionless form, allowing their relative magnitudes to be evaluated. The resulting solution can then be redimensionalized using the same scales to provide an engineering solution in physical units. Here, the horizontal distance is scaled with the characteristic roll-gap length, $\hat{\ell}$, so that $x=\hat{x}$$\hat{\ell}$ gives a dimensionless horizontal coordinate $x$, and the dimensional horizontal position $\hat{x}$ can be recovered from $x$ using $\hat{x}=\hat{\ell}x$. Similarly, the vertical distance is scaled with the initial sheet half-thickness, $\hat{h}_{\!\>\!\mathrm{A}}$, so that $y=\hat{y}/\hat{h}_{\!\>\!\mathrm{A}}$. For circular rolls of radius $\hat{R}$, the roll-gap length $\hat{\ell}$ is equal to
$\hat{\ell}=\sqrt{2\hat{R}(\hat{h}_{\!\>\!\mathrm{A}}-\hat{h}_{1})-(\hat{h}_{\!\>\!\mathrm{A}}-\hat{h}_{1})^{2}},$ (7)
@@ 2.2 Scalings and the non-dimensionalised governing equations
[…]
@@
Hydrostatic pressure, $\hat{p}$, and the normal Cauchy stresses, $\hat{\sigma}_{xx}$ and $\hat{\sigma}_{yy}$, and their deviatoric components, $\hat{s}_{xx}$ and $\hat{s}_{yy}$, are all scaled with the shear yield stress, $\hat{\kappa}$. The assumption of a small friction coefficient is encoded by setting $\beta=\mu/\delta$, where $\beta$ is a quantity of approximate unit magnitude. Similarly, the shear stress, $\hat{\sigma}_{xy}$ is scaled with $\delta\hat{\kappa}$ resulting from small friction. To determine the appropriate order of magnitude of stress components in the $z$-direction, we analysed the boundary condition at the lateral edges of the wire. Since these edges are free surfaces, the material is unconstrained in the $z$-direction and must satisfy the traction-free boundary condition
$\boldsymbol{\hat{\sigma}\cdot m}=0,$ (8)
where $\boldsymbol{m}$ is the unit vector normal to the wire lateral edges. By enforcing this physical constraint, we can infer the scaling of $z$-direction stress components relative to $\hat{\kappa}$, and thus determine their order of magnitude in the model. For a rectangular cross section, the wire lateral edges are given by $\hat{z}-\hat{b}(\hat{x})=0$, so the unit vector normal $\boldsymbol{m}$ is given by
$\boldsymbol{m}=\frac{\hat{\boldsymbol{\nabla}}\big(\hat{z}-\hat{b}(\hat{x})\big)}{\big\|\hat{\boldsymbol{\nabla}}\big(\hat{z}-\hat{b}(\hat{x})\big)\big\|}=\frac{\big(-\delta\mathrm{d}b/\mathrm{d}x,0,1\big)}{\sqrt{1+\delta^{2}\big(\mathrm{d}b/\mathrm{d}x\big)^{2}}}$ (9)
[…]
$\delta\frac{\mathrm{d}b}{\mathrm{d}x}\sigma_{xx}$ $=\hat{\sigma}_{xz}/\hat{\kappa}$ $\text{at }z=b(x),$ (10a)
$\delta^{2}\frac{\mathrm{d}b}{\mathrm{d}x}\sigma_{xy}$ $=\hat{\sigma}_{yz}/\hat{\kappa}$ $\text{at }z=b(x),$ (10b)
$\delta\frac{\mathrm{d}b}{\mathrm{d}x}\hat{\sigma}_{xz}/\hat{\kappa}$ $=\hat{\sigma}_{zz}/\hat{\kappa}$ $\text{at }z=b(x),$ (10c)
[…]
$\hat{x}$ $=\hat{\ell}x,$ $\hat{y}$ $=\hat{h}_{\!\>\!\mathrm{A}}y,$ $\hat{z}$ $=\hat{h}_{\!\>\!\mathrm{A}}z,$
$\hat{h}$ $=\hat{h}_{\!\>\!\mathrm{A}}h,$ $\hat{b}$ $=\hat{h}_{\!\>\!\mathrm{A}}b,$ $\hat{h}_{\!\>\!\mathrm{A}}$ $=\delta\hat{\ell}$
$\hat{\sigma}_{xx}$ $=\hat{\kappa}\sigma_{xx},$ $\hat{\sigma}_{yy}$ $=\hat{\kappa}\sigma_{yy},$ $\hat{\sigma}_{zz}$ $=\delta^{2}\hat{\kappa}\sigma_{zz},$
$\hat{\sigma}_{xy}$ $=\delta\hat{\kappa}\sigma_{xy},$ $\hat{\sigma}_{yz}$ $=\delta^{2}\hat{\kappa}\sigma_{yz},$ $\hat{\sigma}_{xz}$ $=\delta\hat{\kappa}\sigma_{xz},$ (11)
$\hat{u}$ $=\hat{U}_{\!\mathrm{A}}u,$ $\hat{v}$ $=\delta\hat{U}_{\!\mathrm{A}}v,$ $\hat{w}$ $=\delta\hat{U}_{\!\mathrm{A}}w,$
$\hat{\lambda}$ $=\frac{\hat{U}_{\!\mathrm{A}}}{\hat{\ell}\hat{\kappa}}\lambda,$ $\hat{p}$ $=\hat{\kappa}p,$ $\mu$ $=\delta\beta.$
We may now use the scalings introduced in equation (11) to write the governing equations in a dimensionless rescaled form. For example, the vertical force balance is written as
$0$ $=\frac{\partial\hat{\sigma}_{xy}}{\partial\hat{x}}+\frac{\partial\hat{\sigma}_{yy}}{\partial\hat{y}}+\frac{\partial\hat{\sigma}_{yz}}{\partial\hat{z}}$
$=\frac{\partial(\delta\hat{\kappa}\sigma_{xy})}{\partial(\hat{\ell}x)}+\frac{\partial(\hat{\kappa}\sigma_{yy})}{\partial(\hat{h}_{\!\>\!\mathrm{A}}y)}+\frac{\partial(\delta^{2}\hat{\kappa}\sigma_{yz})}{\partial(\hat{h}_{\!\>\!\mathrm{A}}z)}$
$=\frac{\delta\hat{\kappa}}{\hat{\ell}}\frac{\partial\sigma_{xy}}{\partial x}+\frac{\hat{\kappa}}{\hat{h}_{\!\>\!\mathrm{A}}}\frac{\partial\sigma_{yy}}{\partial y}+\frac{\delta^{2}\hat{\kappa}}{\hat{h}_{\!\>\!\mathrm{A}}}\frac{\partial\sigma_{yz}}{\partial z}$
@@
$=\frac{\hat{\kappa}}{\hat{h}_{\!\>\!\mathrm{A}}}\!\left(\!\delta^{2}\frac{\partial\sigma_{xy}}{\partial x}+\frac{\partial\sigma_{yy}}{\partial y}+\delta^{2}\frac{\partial\sigma_{yz}}{\partial z}\right),$ (12)
where $\delta=\hat{h}_{\!\>\!\mathrm{A}}/\hat{\ell}$ is used in the last line. Doing this to the force balance in each direction gives
@@
We now observe the emergence of $\delta$ in the nondimensional equations (13) arising from the scaling defined in (11). This is what enables the asymptotic analysis that follows: for example, from (13b) we see that $\partial\sigma_{yy}/\partial y=O(\delta^{2})$, and so the through thickness variation in $\sigma_{yy}$ is small, and is quantified to be of size $O(\delta^{2}$). This would not have been evident without nondimensionalizing. Similarly, using the scaling in equation (11) the dimensionless forms of the incompressibility and flow rule relations are
$\frac{\partial{u}}{\partial{x}}+\frac{\partial{v}}{\partial{y}}+\frac{\partial{w}}{\partial{z}}$ $=0,$ (14a)
@@
$\frac{\partial{u}}{\partial{y}}+\delta^{2}\frac{\partial{v}}{\partial{x}}$ $=2\delta^{2}\lambda\sigma_{xy},$ (14e)
$\frac{\partial{u}}{\partial{z}}+\delta^{2}\frac{\partial{w}}{\partial{x}}$ $=2\delta^{2}\lambda\sigma_{xz},$ (14f)
$\frac{\partial{v}}{\partial{z}}+\frac{\partial{w}}{\partial{y}}$ $=2\delta^{2}\lambda\sigma_{yz}.$ (14g)